20 (1 Point The null space of A is equal to the number of independent vectors in the rows of the matrix A. true false
Answer #1

The statement is false. ena Take - མ 2 3 O 0 Here one linearly independent row But N(A) =3-R(A) = 3-1 = 2 #1.